[Ncep.nhc.nco_contacts] HSU Shift Log -- Technical Edition -- Sat Sep 23, 2017

HSU Shift Log nhc.tsbadmin at noaa.gov
Sat Sep 23 20:50:52 UTC 2017


HSU Shift Log for Sat Sep 23 2017 16:50:45 GMT-0400 (EDT)

Shift: H

Forecaster: R Berg

Technical Issues Summary:

ATCF issues:

N-AWIPS issues:

AWIPS-2/NCP issues:

Other Techincal issues: Cable is not working in HSU, even though it's  
working in TAFB and the media room

Miscellaneous issues:
-------------- next part --------------
An HTML attachment was scrubbed...
URL: https://www.lstsrv.ncep.noaa.gov/pipermail/ncep.nhc.nco_contacts/attachments/20170923/d8a6656d/attachment.html 


More information about the Ncep.nhc.nco_contacts mailing list